# CLEAR: Character Unlearning in Textual and Visual Modalities

### v1

@@ 4.2 Splits
Ultimately, we utilize the following four splits (sets) to evaluate unlearning:
Forget. Following methodology from Maini et al. (2024), $D_{F}$ is made from data of 2, 10, and 20 persons (1%, 5% and 10% correspondingly) of the full set $D$, consisting of 200 authors. This $D_{F}$ is expected to be unlearned by the model.
Retain. Retain set $D_{R}$ is made from all the other data from the complete set $D$, not included in the $D_{F}$. The model should continue to work well on this subset and preserve its performance as much as possible.
@@
Figure 1 presents a random sample from each of these splits.
@@ 4.3 Evaluation Metrics
To comprehensively evaluate unlearning performance on textual, visual, and textual-visual domains, we measure the MU and MMU performance in terms of the following metrics:
ROUGE-L We calculate the ROUGE-L Lin (2004) score between the model’s predictions and ground truth answers. This metric measures how much the model remembers in exact formulations. However, the generation of models does not always represent the inner knowledge of one’s. That is where the next metric is used.
@@
Probability Score One way of exposing elicit knowledge from a model is through its logits, which are assigned to some factual tokens. We define the conditional probability $p(y|x)^{\frac{1}{|y|}}$ for input $x$ and answer $y$ (power $\frac{1}{|y|}$ corresponds to normalizing for length). Each input question $x$ is considered as a multiple choice question with possible answers $y_{1},...,y_{n}$, and then, assuming $y_{1}$ is the correct answer, desired probability score is computed as $\frac{p(y_{1}|x)}{\sum\limits_{i=1}^{n}p(y_{i}|x)}$. It will be bounded between 0 and 1. A lower probability score indicates that the model is less confident about generating content.
@@
Truth Ratio quantifies the alignment between predictions and the ground truth by comparing the probability of a paraphrased correct answer against the averaged probabilities of several similarly formatted incorrect answers, providing insight into the effectiveness of the unlearning algorithm in removing specific information while maintaining overall accuracy. Assume that $\hat{y}$ denotes a paraphrased version of the answer $y$ for input $x$ and $Y^{\prime}$ is the set of 5 perturbations of the answer $y$. Then desired truth ratio $R$ is calculated as:
$R=\frac{\frac{1}{|Y^{\prime}|}\sum\limits_{y^{\prime}\in Y^{\prime}}p(y^{\prime}|x)^{\frac{1}{|y^{\prime}|}}}{p(\hat{y}|x)^{\frac{1}{|\hat{y}|}}}.$ (3)
This ratio is normalized and rescaled between 0 and 1, with higher values indicating improved unlearning.
Forget Quality. Measuring the quality of forgetting in MU presents significant challenges. The objective is to create a model that cannot be distinguished from one trained solely on retain. The established Hayes et al. (2024) way of measuring the unlearning quality is calculating the U-LIRA score. However, it requires training at least 128 model copies, which is computationally expensive for LLM. A feasible method for achieving this is proposed. We calculate a statistical test on the outputs of two models: our unlearned model and the gold model. The Truth Ratio metric is considered as output for its effectiveness in informativeness. To assess this metric, the Kolmogorov-Smirnov test (KS-Test) is employed to compare the distributions of Truth Ratios from both models. A high p-value from this test suggests effective forgetting, while a low p-value indicates potential privacy leakage and poor unlearning. We call this p-value the Forget Quality of the unlearning method.
Also, we define Real, Retain, and Forget metrics as a harmonical mean of ROUGE, Real Probability score, and Truth Ratio.
@@
LLama2-7B Retain FT 0.50 0.26 0.42 -4.92

### v3

@@ 4.2 Splits
We utilize four splits (sets) to evaluate MU (see Fig. 3 for a sample from these splits):
Forget. Following methodology from Maini et al. (2024), $D_{F}$ is made from data of 2, 10, and 20 persons (1%, 5% and 10%, respectively) of the full set $D$, consisting of 200 authors. This $D_{F}$ is expected to be unlearned by the model.
Retain. The retain set $D_{R}$ consists of all data from the complete set $D$ that is not in $D_{F}$. The model should continue to work well on this subset and preserve its performance as much as possible.
@@
5 Experimental Setup and Evaluation
In this section, we briefly discuss the evaluation metrics and implementation details.
@@ 5.1 Evaluation Metrics
We conduct a comprehensive evaluation using ROUGE-L, Probability Score, Truth Ratio, and Forget Quality metrics to thoroughly assess unlearning performance across textual, visual, and multimodal domains. Following Maini et al. (2024); Li et al. (2024a); Xing et al. (2024), this evaluation setup ensures that we capture the effectiveness of unlearning algorithms while examining both retention and forgetfulness of information within models.
ROUGE-L.
ROUGE evaluates the word-level correspondence between the model’s output and the ground truth answer to a question. We calculate the ROUGE-L recall score Lin (2004) by comparing the model’s decoded output $f_{\hat{\theta}}(x)$ with the ground truth answer $y$ of the gold model $g$: $\text{ROUGE}(f_{\hat{\theta}}(x),y)$. This metric measures the model’s remembrance of the knowledge in its exact formulations.
@@
Probability Score. One way to expose implicit knowledge from a model is through its logits, which are assigned to some factual tokens. This metric assesses the model’s capability to generate the correct answer. We define the conditional probability $p(y|x)^{\frac{1}{|y|}}$ for input $x$ and answer $y$ (power $\frac{1}{|y|}$ corresponds to normalizing for length). Each input question $x$ is considered as a multiple choice question with possible answers $y_{1},...,y_{n}$, and then, assuming that $y_{1}$ is the correct answer, the desired probability score is computed as ${p(y_{1}|x)}/\left(\sum\limits_{i=1}^{n}p(y_{i}|x)\right)$. Higher values indicate better performance, revealing how well the model retains the correct answer.
@@
Truth Ratio quantifies the alignment between predictions and the ground truth by comparing the probability of a paraphrased correct answer against the averaged probabilities of several similarly formatted incorrect answers, providing insight into the effectiveness of the unlearning algorithm in removing specific information while maintaining overall accuracy. As defined by Maini et al. (2024), assume that $\hat{y}$ denotes a paraphrased version of the answer $y$ for the input $x$ and $Y^{\prime}$ is the set of 5 perturbations of the answer $y$. Then desired truth ratio $R$ is calculated as: $R=\frac{1}{\left|Y^{\prime}\right|}\left(\sum_{y^{\prime}\in Y^{\prime}}p\left(y^{\prime}\mid x\right)^{1/\left|y^{\prime}\right|}\right)/p(\hat{y}\mid x)^{1/|\hat{y}|}$. This ratio is normalized and rescaled between 0 and 1, with higher values indicating better knowledge retention.
Aggregate metrics. All three above-defined metrics are bounded between 0 and 1, so we combine them into a single metric to evaluate the overall performance. We set up the Real, Retain, and Forget metrics as a harmonic mean of the ROUGE, the Probability score, and the Truth Ratio computed on corresponding dataset splits.
Forget Quality calculates the "distance" of the unlearned model to the gold model, which is a proxy metric for the quality of exact unlearning. Following Maini et al. (2024), we take the Truth Ratios distribution of both models. But, instead of the p-value of the Kolmogorov-Smirnov test, we calculate the Jensen-Shannonn distance between these distributions. The latter metric better captures the differences between models, which we additionally check and describe in the Appx. E. To maintain the higher - the better convention, we subtract the distance from 1.
Method Text-only Visual-only Multimodal
(Forget/Retain) (Forget/Retain) (Forget/Retain)
LLMU 0.01/0.03 85.2/88.9 0.25/0.51
DPO 0.42/0.26 50.2/81.4 0.22/0.48
SCRUB 0.42/0.26 42.59/99.4 0.36/0.52
IDK 0.24/0.26 N/A 0.33/0.51
RMU 0.59/0.26 67.9/99.0 0.00/0.00
Retain FT 0.42/0.26 100.0/100.0 0.37/0.51
Performance correlation with multimodal:
Spearman’s $\rho$ 0.705 (p=0.118) 0.205 (p=0.741) 1.00
Table 2: Transferability analysis across domains. We report forget (F) and retain (R) metrics for each method. Lower F and higher R are better. N/A indicates the method was not applicable. Correlation shows Spearman’s rank correlation between single-domain and multimodal performance, with p-values in parentheses.
5.2 Implementation
For the source model, we use LLaVa model Liu et al. (2023) with ViT Dosovitskiy et al. (2021) as visual encoder and LLaMa2-7B Touvron et al. (2023) as a language model. First, we finetune it on the image captioning task using the full CLEAR, both visual and textual parts. We call this model “base”, as it contains forget and retain sets. Then, we perform the unlearning process on it. We use the same hyperparameters for each method. We evaluate the unlearned model based on our metrics from Sec. 5, using the Multi-choice VQA task for the probability score and the image captioning task for the Truth Ratio. For comparison, we also present the metrics of the “gold” model. Experiment results and metrics are shown in Tab. 1, with details provided in Appx. H.
6 Results
Method Modality Real $\uparrow$ metric Retain $\uparrow$ metric Forget $\downarrow$ metric Forget $\uparrow$ Quality
Gold 0.50 0.51 0.19 1.00
Base 0.48 0.51 0.35 0.85
RMU text 0.31 0.01 0.02 0.75
RMU visual 0.24 0.00 0.00 0.75
RMU both 0.22 0.00 0.00 0.80
GD text 0.26 0.00 0.00 0.79
GD visual 0.29 0.00 0.00 0.20
GD both 0.49 0.51 0.37 0.85
Retain FT text 0.49 0.51 0.37 0.85
Retain FT visual 0.46 0.45 0.42 0.85
Retain FT both 0.46 0.46 0.40 0.85
GA text 0.34 0.00 0.00 0.22
GA visual 0.29 0.00 0.00 0.32
GA both 0.29 0.00 0.00 0.32
KL text 0.49 0.32 0.23 0.71
KL visual 0.29 0.00 0.00 0.28
KL both 0.48 0.35 0.27 0.81
IDK text 0.48 0.50 0.39 0.85
IDK visual 0.44 0.45 0.35 0.84
IDK both 0.46 0.46 0.33 0.84
NPO text 0.51 0.51 0.29 0.85
NPO visual 0.48 0.43 0.24 0.84
NPO both 0.49 0.00 0.00 0.72
SCRUB text 0.49 0.51 0.37 0.85
SCRUB visual 0.48 0.49 0.39 0.85
SCRUB both 0.49 0.51 0.37 0.85
DPO text 0.47 0.45 0.42 0.85
DPO visual 0.48 0.49 0.23 0.84
DPO both 0.46 0.47 0.28 0.84
LLMU text 0.48 0.46 0.40 0.85
LLMU visual 0.49 0.50 0.37 0.85
LLMU both 0.47 0.48 0.33 0.84
Table 3: Results of unlearning of different modalities within multimodal setup. We finetune on full datasets (both modalities), then forget on a single domain subset (text or visual) or full forget set. Base – model before unlearning. Gold - a model trained only on retain.
In the following sections, we describe the results of MU methods on our dataset. We seek to answer the following research questions:
  1. RQ1:
    Does an unlearning method’s performance on a single domain transfer directly to the performance in a multimodal setting? Should we study multimodal unlearning at all if we can easily predict its performance from single-domain experiments?
  2. RQ2:
    Is unlearning only one modality (textual or visual) enough in a multimodal setup? How does the effectiveness vary depending on modality? How do different unlearning methods compare in their effectiveness for unlearning specific modalities?
  3. RQ3:
    In the context of multimodal unlearning, what methods perform the best?
6.1 Transferability from Single Domain
To investigate RQ1, we analyse how well the performance of unlearning methods in single modalities predicts their effectiveness in multimodal settings. For the textual domain, we use the TOFU benchmark; for the visual, we use a standard U-MIA approach to the data, consisting of the faces from our dataset; the details for the pipelines and full results are provided in Appx. F and G. For multimodal unlearning, we use our benchmark.
The correlations between single-domain and multimodal (MM) rankings are relatively weak. We rank methods according to their forget metric performance in each domain and calculate Spearman’s rank correlation coefficient between single-domain and multimodal rankings. The correlation $\rho=0.7$ for text-MM and $\rho=0.2$ for visual-MM, indicating limited transferability (see Tab. 2).
We observe significant discrepancies between single-domain and multimodal performance. For example, LLMU achieves a retain metric of 0.51 in multimodal but degrades to 0.03 in the textual setting while maintaining good forget scores (0.25 vs 0.01). Similar patterns emerge for other methods, suggesting that single-domain evaluation is insufficient.
Methods that perform well in single domains can fail catastrophically in multimodal settings. For instance, RMU achieves good forget-retain balance in text-only (0.26/0.59) but completely fails on the retain set in multimodal setup (0.00/0.00), highlighting the unique challenges of multimodal unlearning.
Takeaway 1: Single-domain performance is a poor predictor of multimodal unlearning success, with relatively low-rank correlations ($\rho=0.7$ and $\rho=0.2$) and distinct failure modes. This emphasizes the need for dedicated multimodal evaluation frameworks and potentially new methods designed specifically for multimodal MU.
6.2 Impact of Modality Selection on Unlearning
To explore RQ2, we examine how the choice of unlearning modality impacts performance. We conduct experiments with three variants for each method: text-only, visual-only, and both modalities. Results in Tab. 3 show distinct patterns across methods.
Text-only Unlearning.
Text-only approaches show mixed results. While some methods, like NPO, achieve good retain metrics (0.51) with moderate forget scores (0.29), others struggle significantly. RMU and GD completely fail on retain (0.01 and 0.00). KL shows middling performance with retain at 0.32 and forget at 0.23. This inconsistency suggests that text-only unlearning may disrupt cross-modal representations unpredictably.
Visual-only Unlearning.
Visual-only unlearning often achieves a better balance. DPO shows promising results with a forget metric of 0.23 while maintaining a 0.49 retain score. LLMU and SCRUB demonstrate similar patterns (forget: 0.37, 0.39; retain: 0.50, 0.49, respectively). However, some methods like RMU, GD, GA, and KL completely fail on retain metrics (0.00), indicating visual-only approaches are not universally successful.
Multimodal Unlearning.
Joint modality unlearning shows the most promising results for several methods. IDK improves its forget metric from 0.39 (text) and 0.35 (visual) to 0.33 (both) while maintaining stable retain performance (0.46). NPO shows strong real-world performance (0.49) but struggles with retain metrics when using both modalities. SCRUB demonstrates remarkable consistency across configurations (retain:  0.50, forget:  0.37), suggesting some methods are more robust to modality selection.
Takeaway 2: While visual-only unlearning often outperforms text-only approaches, the effectiveness varies significantly by method. Methods like SCRUB maintain consistent performance across modalities (retain: 0.49-0.51), while others show dramatic variations. NPO and KL demonstrate that combining modalities can improve forget quality (0.72-0.81) compared to single-modality approaches (0.28-0.85). However, the optimal choice of modality depends heavily on the specific method and desired performance trade-offs.
6.3 Unlearning Both Domains
Having established that multimodal unlearning requires addressing both modalities, we evaluate all available unlearning methods on our source model $f_{\theta}$ across both domains. For these experiments, we use a forget set containing data about 20 persons (10% of the dataset), encompassing both their textual and visual information.
As shown in Tab. 3, there are three distinct categories of method behaviour. GA, GD, KL, and RMU achieve perfect unlearning (forget metric = 0) but completely destroy the model’s retained knowledge (retain metric = 0). IDK, SCRUB, and Retain FT maintain strong retain performance ( 0.51) but struggle with effective forgetting (forget metrics 0.33-0.37). LLMU and DPO balance forgetting and retention best, maintaining reasonable retain metrics (0.48-0.51) while showing improved forget performance (0.22-0.25).
Takeaway 3: Most unlearning methods struggle with the trade-off between effective forgetting and knowledge retention in multimodal settings. Only LLMU and DPO show promise in balancing these objectives, but their performance remains below the gold model (forget = 0.19, retain = 0.51).
7 Conclusion
In this work, we introduce CLEAR, the first open-sourced benchmark designed to assess machine unlearning in a textual-visual multimodal setup. Our evaluation of existing unlearning techniques across domains shows that multimodal unlearning is more challenging than previously anticipated, laying the ground for further research. Our findings offer a new perspective than earlier results on safety alignment Chakraborty et al. (2024), which suggested that text-only unlearning is sufficient for multimodal models.
While CLEAR’s synthetic personas ensure controlled evaluation, real-world data (e.g., diverse facial features, noisy captions) may introduce new challenges. Additionally, our study focuses on visual-language models, leaving other modalities (e.g., audio, video) unexplored. By open-sourcing CLEAR and establishing the first multimodal MU leaderboard, we aim to accelerate progress toward ethical, privacy-preserving multimodal AI. Our findings highlight that MMU is not merely an extension of unimodal unlearning but a distinct challenge requiring novel methodologies.
Limitations
Ethics
In this work, we focus on unlearning character-specific knowledge in pre-trained visual-language models (VLMs). We aim to enable VLMs to selectively forget all traces of specific synthetic personas—including their textual biographies, visual appearances, and cross-modal associations—while preserving the model’s general capabilities. This addresses critical ethical concerns, such as the right to be forgotten and prevention of unintended memorization. For forget and retain sets, all data is synthetically generated to avoid biases and leakage from real-world sources, with evaluation protocols encouraging responsible use. These sets were manually checked by one of the authors. Datasets on celebrity recognition and general VQA are publicly accessible sources. We also urge researchers and developers to employ our methods responsibly and with ethical considerations.
We used 84 hours of A100 GPU computation, resulting in an estimated 9 kg of CO2 emissions.
Use of AI Assistants.
We utilize Grammarly to enhance and proofread the text of this paper, correct grammatical, spelling, and stylistic errors, as well as rephrase sentences. Consequently, certain sections of our publication may be identified as AI-generated, AI-edited, or a combination of human and AI contributions.
Appendix A Rationale for Synthetic Face Selection
The use of synthetic faces, rather than real-world facial data, in our benchmark is motivated by ethical, technical, and practical considerations. First, synthetic faces eliminate privacy risks and ethical concerns associated with real facial datasets. By generating artificial personas, we avoid biases inherent in real-world datasets and ensure no real individuals are misrepresented, aligning with the right to be forgotten principle.
Second, synthetic data provides precise control over memorization evaluation. Real faces risk contamination from prior model training (e.g., pre-existing celebrity images in model weights), which could confound unlearning performance measurements. Synthetic faces, being novel and never publicly released, guarantee that models learn exclusively from our benchmark, enabling accurate assessment of unlearning efficacy. Notably, our experiments reveal that models still struggle to fully erase synthetic faces – despite their controlled generation. This implies that applying current unlearning methods to real-world faces (e.g., from public sources like Wikipedia) would face greater challenges, as real data introduces uncontrolled variability and pre-existing biases that synthetic benchmarks deliberately exclude. In other words, applying artificial profiles ensures that the considered model has not seen the authors during pre-training, and this is essential for the fair evaluation of MU methods, as we can easily compare MU results with a gold model, which has never seen the profiles we want to forget, without expensive re-training from scratch on a large plethora of data required for LLMs and VLLMs training.
Third, synthetic generation prevents cross-modal leakage. By explicitly linking synthetic faces to their textual biographies, we isolate memorization tests to our dataset, ensuring no external knowledge interferes. This allows rigorous evaluation of whether unlearning a biography also removes its associated face.
Additionally, synthetic faces enhance reproducibility and scalability. Unlike real datasets burdened by licensing restrictions, synthetic data can be freely shared, fostering open benchmarking. On-demand generation also supports customizable testing, such as expanding the forget set to thousands of unique identities without legal barriers. Our comprehensive image generation strategy suits the author’s textual descriptions and preserves consistency among his or her images. Still, it enables sufficient diversity between different authors regarding age, gender and ethnicity.
Also, it is worth noting that synthetic data is also used in practice in literature for unlearning-related task. In Zhang et al. (2024b), the authors introduce a novel dataset, UnlearnCanvas, designed to benchmark machine unlearning techniques in diffusion models, offering a comprehensive, high-resolution stylized image dataset to evaluate the unlearning of artistic styles and associated objects. The UnlearnCanvas dataset includes generated images across 60 artistic painting styles, with 400 images per style across 20 object categories. The dataset facilitates the quantitative evaluation of vision generative modelling tasks, including machine unlearning, style transfer, vision in-context learning, bias removal for generative models, and out-of-distribution learning. The paper Ma et al. (2024) introduces a new benchmark, FIUBench, to evaluate the effectiveness of unlearning algorithms in Vision Language Models under the Right to be Forgotten setting. The authors formalize the VLM unlearning task and construct a Fictitious Facial Identity VQA dataset of synthetic faces paired with randomly generated personal information to study privacy under the Right to be Forgotten scenario. This approach allows precise control over the source of information and its exposure in the unlearning dataset. The dataset includes personal backgrounds, health records, and criminal histories for each facial identity. The work Dhasade et al. (2024) introduces a novel approach to Federated Unlearning, which aims to effectively remove specific training data knowledge from machine learning models trained through Federated Learning. The authors highlight the inefficiencies of existing Federated Unlearning methods that often involve high computational costs due to gradient recomputation and storage requirements. The provided approach, QuickDrop, is designed to streamline the unlearning process by generating compact synthetic datasets that represent the gradient information used during model training. This approach significantly reduces the volume of data needed for unlearning while maintaining performance efficiency. QuickDrop employs a method called dataset distillation to create a compact dataset that captures essential features of the original training data. This dataset is approximately 1% of the size of the original data, leading to minimal storage overhead. Each client generates a synthetic dataset through gradient matching, which serves as a compressed representation of their original gradients.
In summary, synthetic faces prioritize ethical rigour, experimental precision, and reproducibility—critical for advancing multimodal machine unlearning research. The observed difficulty in unlearning even synthetic faces underscores fundamental model limitations, which real-world deployments (e.g., authors’ faces) would exacerbate due to added complexity. Our benchmark thus serves as a necessary precursor to addressing practical challenges in ethical AI.
Appendix B Unlearning Methods
  1. 1.
  2. 2.
    $L(D_{F},\theta)=\frac{1}{|D_{F}|}\sum_{x\in D_{F}}{NLL}(x,\theta),$
  3. 3.
    $L_{GD}=-L(D_{F},\theta)+L(D_{R},\theta),$
  4. 4.
    The total objective function is formulated as follows:
    $L_{obj}=-L(D_{F},\theta)+L_{\text{KL}},$
  5. 5.
    $L_{idk}=L(D_{R},\theta)+L(D_{F}^{idk},\theta),$
  6. 6.
  7. 7.
    Negative Preference Optimization . Proposed in Zhang et al. (2024a) this method can be treated as DPO without positive examples. In our setting, the final loss function $L_{NPO}$ for this method is derived as follows:
  8. 8.
    $d(x,w^{s})=\operatorname{KL}(p(f(x;w^{o}))||p(f(x;w^{s}))),$
    $L_{R}=\frac{\alpha}{|D_{R}|}\sum_{x_{r}\in D_{R}}d(x_{r},w^{s}),$
    $L_{F}=\frac{1}{|D_{F}|}\sum_{x_{f}\in D_{F}}d(x_{f},w^{s}),$
    $L_{\text{task}}=\frac{\gamma}{|D_{R}|}\sum_{x_{r}\in D_{R}}l(x_{r},y_{r}),$
    $L=L_{R}-L_{F}+L_{\text{task}},$
  9. 9.
    LLMU Yao et al. (2024c)
    This method was proposed in one of the first works on unlearning LLMs Yao et al. (2024c). In our experiments, we made slight modifications to the original method, and employed the following loss function:
    $L_{F}:=-L(D_{F},\theta),$
    $L_{LLMU}=L_{F}+L_{r}+L_{R},$
  10. 10.
    $L_{\operatorname{RMU}}=L_{\operatorname{F}}+L_{\operatorname{R}},$
  11. 11.
    $L_{\operatorname{F}}=d(f(x_{f}),f_{o}(x_{f}^{aug})),$
    $L_{\operatorname{R}}=d(f(x_{r}),f_{o}(x_{r})),$
    $L=L_{\operatorname{F}}+L_{r},$
  12. 12.
    $L=L_{\operatorname{F}}+L_{R},$
  13. 13.
    $L=L_{\operatorname{R}}+\lambda\cdot||\theta||_{1},$
    where $\lambda$ is a parameter of regularization.
  14. 14.
    Selective Knowledge Unlearning. Liu et al. (2024). This method is based on the weights arithmetic. First, we additionally finetune the model on the forget set with this loss:
    $L_{GD}=\sum_{(x_{f},y_{f})\in D_{F}}l(f(x_{f}),y_{f})$
    $L_{RD}=\sum_{x^{i}_{f}\in D_{F}}\frac{1}{Y^{i}_{rd}}\sum_{y\in Y^{i}_{rd}}l(f(x^{i}_{f}),y)$
    $L_{PD}=\sum_{(x_{r},y_{r}}KL(p(x_{r}),y_{r})$
    $L=\epsilon_{1}\cdot L_{GD}+\epsilon_{2}\cdot L_{RD}-\epsilon_{3}\cdot L_{PD}$
    Where $Y^{i}_{rd}$ is the set of related answers to the given question $x_{i}$. So, the finetuned version is the opposite of what we aim to achieve. Then, we calculate the delta in weights, produced by this finetuning, and substract it from the original model.
Appendix C The process of face generation
@@
LLama2-7B Original 0.47 0.26 0.42 -3.92
Gold 0.48 0.26 0.24 0.0
Retain FT 0.50 0.26 0.42 -4.92